Let the function f be defined by f(x)=x^2-7x+10 and f(t+1)=0, what is one possible value of t?
Answer by bloggauta(8)

f(x)= x^2-7x+10 and f(t+1)= 0
f(t+1)= (t+1)^2-7(t+1)+10= 0
t^2+2t+1-7t-7+10= 0
t^2-5t+4= 0 → (t-4)(t-1)= 0
t= 4 or t= 1